OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (1, -9) and (3, -1). Please show all of your work

OpenStudy (anonymous):

(y-y1)/(x-x1) = m so, slope: m=(-1-(-9))/(3-1)=4 y-(-1)=4*(x-3) y+1=4x-12 y=4x-13